OpenStudy (firejay5):

Geometry. Show work and explain. 23. A rectangular pyramid has a base area of x^2 + 3x - 10 over 2x square centimeters and a height of x^2 - 3x over x^2 - 5x + 6 centimeters. Write a rational expression to describe the volume of the rectangular pyramid.

OpenStudy (anonymous):

\[A=\frac{1}{2}bh\] \[x^2-4=\frac{1}{2}(2x+4)h\] \[x^2-4=(x+2)h\] and since \[x^2-4=(x+2)(x-2)=(x+2)h\] you know \(h=x-2\)

OpenStudy (mertsj):

The volume of a pyramid is 1/3 Bh were B is the area of the base.

OpenStudy (mertsj):

\[V=\frac{1}{3}\times \frac{x^2+3x-10}{2x} \times \frac{x^2-3x}{x^2-5x+6}\]

OpenStudy (mertsj):

I'm assuming that you would factor and simplify that.

OpenStudy (firejay5):

I factored, what would I do with the 1/3

OpenStudy (mertsj):

Just make 3 a factor of the denominator.

OpenStudy (firejay5):

for my answer could I keep the 1 at top and 3 on bottom

OpenStudy (mertsj):

\[\frac{(x+5)(x-2)}{3(2x)} \times\frac{x(x-3)}{(x-3)(x-2)}=\frac{x+5}{6}\]

OpenStudy (firejay5):

You got a medal thanks! :DDDDDD @Mertsj

OpenStudy (mertsj):

yw